OpenStudy (anonymous):

QUICK HELP? Solve the following system of equations by linear combination: 2d – e = 7 d + e = 5 There is no solution. There are an infinite number of solutions. The solution is (3, –1). The solution is (4, 1).

OpenStudy (amoodarya):

add them adding the equation is a linear combination

OpenStudy (amoodarya):

by adding you will have 2d-e+d+e=7+5 3d=12 d=4 now put d in one of them 2d-e=7 8-e=7 e=1 (d,e)=(4,1)